# Loop Series Expansions for Tensor Networks

### v1

@@ I Introduction
Tensor networks (TNs) are powerful tools for the classical simulation of quantum systems TN1 ; TN2 ; TN3 ; TN4 ; TN5 ; TN6 ; TN7 ; TN8 ; TN9 ; TN10 . Although they were traditionally developed for model systems in the context of condensed matter physics, they have more recently been applied towards the benchmarking of nascent quantum computers QC1 ; QC2 ; QC3 ; QC4 ; QC5 ; QC6 ; QC7 ; QC8 ; QC9 ; QC10 ; QC11 ; QC12 . Tensor network states were pioneered by the introduction of matrix product states MPS1 ; MPS2 (MPS), which possess tensors connected in a $1D$ geometry. For several decades numeric algorithms based on the MPS, such as the density matrix renormalization groupDMRG1 ; DMRG2 ; DMRG3 (DMRG), have been widely regarded as the gold-standard of classical simulation algorithms for $1D$ quantum systems; producing high-accuracy and unbiased results for a wide range of problems. A large part of this success can be attributed to the computational efficiency with which MPS can be contracted. Subsequent developments have seen the proposal of sophisticated tensor networks, such as projected entangled pair statesPEPS1 ; PEPS2 (PEPS) and multi-scale entanglement renormalization ansatzMERA1 ; MERA2 (MERA), designed to efficiently represent quantum states on large $2D$ systems. However, a significant challenge of these $2D$ tensor networks is the high computational expense required for their contraction PEPS3 even when approximate contraction schemes are used PEPS4 ; PEPS5 ; PEPS6 ; PEPS7 ; PEPS8 . In practice, this expense limits the amount of entanglement that can be encoded in these networks and therefore also limits the range of problems for which they produce accurate results. Thus the development of strategies for more efficient contraction of TNs remains a vital task.
A promising approach for the contraction of TNs is through the adoption of belief propagationBP0 ; BP1 ; BP2 (BP), a message passing algorithm with a long history in both computer scienceBP3 ; BP4 ; BP5 and in statistical physicsBP6 ; BP7 ; BP8 . In recent times the application of BP towards TNs has generated much interest BP9 ; BP10 ; BP11 ; BP12 ; BP13 ; BP14 ; BP15 ; BP16 ; BP17 ; BP18 , and produced a notable success QC11 ; QC12 of simulating a quantum experiment performed on IBMs 127-qubit Eagle processor IBM1 , a problem that was seen to be intractable for other $2D$ TN algorithms ISO1 . However, a major drawback of BP is that it is an uncontrolled approximation; one cannot systematically improve the accuracy of a BP result as is otherwise done in the context of tensor networks by simply growing the bond dimension, although there do exist schemes to improve the accuracy of BP through a preliminary clustering of a network BP15 .
@@ II Belief Propagation
which is known as the Bethe free energy BP0 .
@@ III Configurations in the BP Basis
with $\mathbb{I}$ as the $d$-dim identity, see also Fig. 1(a). We henceforth refer to $\mathcal{P}_{rs}$ as the projector onto the (BP) ground state of edge $(r,s)$ and $\mathcal{P}_{rs}^{\textrm{C}}$ as the projector onto the excited subspace. It follows that the tensor network $\mathcal{T}$ can be resolved as a sum over $2^{M}$ (BP basis) configurations, formed from all combinations of projecting each edge into either the ground or excited sub-space. We define the degree of a configuration as the number of excited edges that it contains, using $\delta_{x}$ to denote a degree-$x$ configuration and $W(\delta_{x})$ to denote its weight (i.e. the scalar resulting from its contraction). Thus we can recognise $\delta_{0}$ as the BP vacuum; the unique configuration with all edges projected into their BP ground state. It follows that $W(\delta_{0})$ evaluates to the standard BP result of Eq. 4. Our goal is now to characterise the weights of the remaining configurations, which may be thought of as excitations on top of the BP vacuum, such that the high-weight contributions can then be re-introduced into the approximation.
@@
see also the example shown in Fig. 1(c). This result is further simplified by re-scaling each tensor $T_{r}$ so that its vacuum state contribution, i.e. $\tilde{T}_{r}$ from Eq. 3, is unity which allows the vacuum component of any configuration to be ignored. This observation implies that one only needs to understand the connected excitations in order to fully characterize a tensor network contraction.
@@ IV.1 Free-energy density
Our first step is to evaluate the weights $W(\delta_{x})$ of the lowest degree excitations $\delta_{x}$; in the calculations performed in this manuscript we evaluate the excitations up to degree $x=12$ as shown in Fig. 2(b-e). Each of these weights is given by contracting a small tensor network formed by capping the external indices (i.e. of tensors within the excitation) with their fixed point messages and projecting the internal indices onto their excited subspace with projectors $\mathcal{P}_{rs}^{\textrm{C}}$ as defined in Eq. 6. However, even once these weights $W(\delta_{x})$ are known, it still remains a non-trivial combinatorial task to compute their impact on the free energy density $f$. An approximate counting of the possible combinations of loop excitations, as derived in Appendix B, gives the following result for the free energy correction
@@
with $L(\delta_{x})$ as the number of locations per lattice site that excitation can be placed and $S(\delta_{x})$ as the number of lattice sites that the excitation occupies. For instance, the $\delta_{6}$ excitation shown in Fig. 1(b) has $L(\delta_{6})=1/2$ (since there is one hexagon per two lattice sites) and $S(\delta_{6})=6$. Notice that the terms $\exp(S(\delta_{x})f)$ in Eq. 11 act to suppress excitations based on the amount of space they occupy. Given that Eq. 11 contains the unknown $f$ on both sides we cannot solve for $f$ directly. Instead we solve Eq. 11 iteratively by first setting $f=0$ on the r.h.s, computing a new value for $f$, then feeding the new $f$ back into r.h.s and repeating until $f$ is converged. Usually this requires only a few iterations to converge to high precision.
@@ IV.2 Transfer matrix
We now examine how one can compute the transfer matrix $\textrm{T}_{AB}$ associated to cutting open a single edge from the network $\mathcal{T}=\mbox{$\langle\psi|\psi\rangle$}$ as shown in Fig. 3(a). Such transfer matrices are utilized in determining the optimal truncation of an internal index from the PEPS and thus play an important role in many PEPS optimization algorithms PEPS4 ; PEPS5 ; PEPS6 ; PEPS7 .
The BP vacuum contribution to $\textrm{T}_{AB}$ is shown in Fig. 3(b) together in Fig. 3(c) with examples of some of the contributions from loop excitations, all of which individually evaluate to a matrix with the same dimensions as $\textrm{T}_{AB}$. The series approximation to $\textrm{T}_{AB}$ is constructed by evaluating each of the loop-excitation networks (up to some desired degree), then adding the contributions together after weighting each by a suppression factor $\exp(S(\delta_{x})f)$ similar to Eq. 11. Note this requires that the free energy density $f$ be evaluated prior to computing $\textrm{T}_{AB}$. Finally, we remark that excitation terms that do not include either of the tensors adjacent to an open index do not need to be considered explicitly; their contributions, which are proportionate to the vacuum term, are already accounted for through the suppression factors.
@@ V Benchmark results
We now benchmark our approach, beginning with contraction of a PEPS representing the ground state $|\psi\rangle$ of the hexagonal lattice AKLT modelAKLT1 ; AKLT2 in the thermodynamic limit. The AKLT model is particular useful as a test bed, since $|\psi\rangle$ can be represented exactlyPEPS1 with an iPEPS of bond dimension $m=2$. Following the methodology prescribed in Sect. IV we compute the free energy density $f$, transfer matrix $\textrm{T}_{AB}$, and the density matrix $\rho_{AB}$ for loop expansions containing up to $12^{\textrm{th}}$ degree excitations $\delta_{12}$ shown in Fig. 2. In all cases we compare against evaluations using a boundary MPS contractionPEPS5 whose dimension $\chi$ has been chosen large enough, $\chi\approx 30$, such that any truncation errors are negligible and the results can be considered as numerically exact. Fig. 4 shows that the addition of the loop corrections improve the accuracy of the BP results dramatically, such that they appear to converge towards the exact results as the degree of the expansion is increased. The expansion to $12^{\textrm{th}}$ degree excitations $\delta_{12}$ gave the free energy density $f$ accurate to $\varepsilon=2\times 10^{-7}$, which represents an improvement of four orders of magnitude over the BP result. That the infinite hexagonal tensor network can be characterised so accurately from only the BP vacuum plus a few of the low-degree loop excitations is remarkable. Note that we also compare against results obtained from the exact contraction of the PEPS on a finite $N\times N$ geometry (with periodic boundaries) for $N=\{6,8,10,12\}$. In all cases we see that an $N^{\textrm{th}}$ degree loop expansion exceeds the accuracy of an exact evaluation on a $N\times N$ geometry. This is expected since the $N\times N$ systems would contain degree-$N$ corrections (i.e. that ‘wrap’ the whole system) whereas a degree-$N$ loop expansion is only neglecting excitations of degree larger than $N$. None-the-less, this provides additional confirmation that the loop expansions are properly accounting for all short-range contributions.
@@
Fig. 5(b-c) presents equivalent results for kagome and square lattices. For iPEPS defined on kagome lattices we perform local manipulations to exactly transform the network into a hexagonal geometry (similar to the form of PEPS described in Ref. PESS1, ) prior to applying the series expansion, see Appendix C for details. The square lattice results were obtained by applying the expansion directly to the square lattice, evaluated up to the $\delta_{8}$ excitations as shown in Appendix C. Similar to the hexagonal lattice, the results from the kagome and square lattices also show robust improvements over the BP approximation. However we see that the convergence on the square lattice is slower with respect to increasing the degree of expansion; we attribute this to the fact that the quantity of distinct loop configurations at any given degree grows much faster than on the hexagonal lattice.
Figure 4: Results comparing the accuracy of BP and loop expansion up to degree $\delta_{x}$ against exact results for the ground state of the (thermodynamic limit) hexagonal AKLT model. Displayed are errors in (a) the free-energy density $f$, (b) the transfer matrix $\textrm{T}_{AB}$, (c) the two-site density matrix $\rho_{AB}$. Dashed lines display the errors given from an exact contraction of an $N\times N$ system.
@@ VI Discussion
Another direct application of these results could be for PEPS optimization algorithms. Given that a recent workBP13 established the equivalence of BP to the ‘simple update’ strategy for PEPS, the series expansion could significantly enhance the accuracy of the ‘simple update’ strategy while still remaining relatively computationally efficient. For instance, the genus-1 loop excitations on the hexagonal lattice, i.e. the $\delta_{6}$ and $\delta_{10}$ terms in Fig. 2, can be evaluated with $O(m^{6})$ cost for PEPS of bond dimension $m$, and this cost could also potentially be reduced via approximation strategies like sparse diagonalization. Thus the dominant terms in the expansion are still feasible to compute even for PEPS with bond dims in excess of $m=100$, well beyond the computational limits of boundary MPS or corner transfer matrix methods for PEPS contractionsPEPS5 ; PEPS6 ; PEPS7 ; PEPS8 .
@@
Most existing tensor network algorithms are built on the twin pillars of tensor contractions and tensor decompositions/truncations. The results presented in this manuscript, which prescribe a general procedure for resolving a complicated network into its fundamental components, could represent a third pillar for tensor network algorithms; one that could have interesting synergies with established routines. For instance, in evaluating a network $\mathcal{T}$, one could begin by expanding in terms of the BP vacuum, the short-loop excitations, and the remaining ‘residual’. Potentially, one could then apply a traditional (approximate) contraction strategy on the residual which, presumably, would be more amenable to truncations due to the separation from the short-loop excitations. Exploring the ways that series expansions could compliment or enhance existing tensor network algorithms remains an interesting avenue for future research.
@@ Acknowledgements.
We thank AWS for supporting the quantum computing program. GKC acknowledges support from the US DOE, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (QSA), and from a generous gift from AWS.
@@ Appendix A Exponential Suppression of High-degree Excitations
The utility of the proposed loop expansions is built on the premise that the weight of excitations (i.e. on top of the BP vacuum) are exponentially diminishing in their degree; this property allows us to accurately approximate a tensor network contraction using only a few low-degree excitations. In this appendix we justify this premise.
@@
Although our argument assumed, for simplicity, that the transfer matrices with any loop were identical this condition is not strictly necessary; it is only required that the transfer matrices have a gap between their dominant and sub-leading eigenvalues in order to realize (approximate) exponential decay. Finding a general characterisation of the networks that are expected to have a gap in the transfer matrix spectrum associated to any closed loop is likely a difficult task (and certainly beyond the scope of this manuscript). However, given that the loop series expansion is only expected to be viable for networks where the loop corrections are small in comparison to the BP approximation, it follows that the gap condition must hold in this circumstance. Finally we remark that our argument is only valid for excitations that realize genus-1 loops not, for instance, higher genus loops like the $\delta_{11}$ excitations shown in Fig. 2(d). None-the-less, it seems to be the case that the scaling of Eq. 15 does hold for higher genus loop excitations.

### v2

@@ I Introduction
Tensor networks (TNs) are powerful tools for the classical simulation of quantum systems TN1 ; TN2 ; TN3 ; TN4 ; TN5 ; TN6 ; TN7 ; TN8 ; TN9 ; TN10 . Although they were traditionally developed for model systems in the context of condensed matter physics, they have more recently been applied towards the benchmarking of nascent quantum computers QC1 ; QC2 ; QC3 ; QC4 ; QC5 ; QC6 ; QC7 ; QC8 ; QC9 ; QC10 ; QC11 ; QC12 . Tensor network states were pioneered by the introduction of matrix product states MPS1 ; MPS2 (MPS), which possess tensors connected in a $1D$ geometry. For several decades numeric algorithms based on the MPS, such as the density matrix renormalization groupDMRG1 ; DMRG2 ; DMRG3 (DMRG), have been widely regarded as the gold-standard of classical simulation algorithms for $1D$ quantum systems; producing high-accuracy and unbiased results for a wide range of problems. A large part of this success can be attributed to the computational efficiency with which MPS can be contracted. Subsequent developments have seen the proposal of sophisticated tensor networks, such as projected entangled pair statesPEPS1 ; PEPS2 (PEPS) and multi-scale entanglement renormalization ansatzMERA1 ; MERA2 (MERA), designed to efficiently represent quantum states on large $2D$ systems. However, a significant challenge of these $2D$ tensor networks is the high computational expense required for their contraction PEPS3 even when approximate contraction schemes are used PEPS4 ; PEPS5 ; PEPS6 ; PEPS7 ; PEPS7b ; PEPS7c ; PEPS8 . In practice, this expense limits the amount of entanglement that can be encoded in these networks and therefore also limits the range of problems for which they produce accurate results. Thus the development of strategies for more efficient contraction of TNs remains a vital task.
A promising approach for the contraction of TNs is through the adoption of belief propagationBP0 ; BP1 ; BP2 (BP), a message passing algorithm with a long history in both computer scienceBP3 ; BP4 ; BP5 and in statistical physicsBP6 ; BP7 ; BP8 . In recent times the application of BP towards TNs has generated much interest BP9 ; BP10 ; BP11 ; BP12 ; BP13 ; BP14 ; BP15 ; BP16 ; BP17 ; BP18 , and produced a notable success QC11 of simulating a quantum experiment performed on IBMs 127-qubit Eagle processor IBM1 , a problem that was seen to be intractable for other $2D$ TN algorithms ISO1 . However, a major drawback of BP is that it is an uncontrolled approximation; one cannot systematically improve the accuracy of a BP result as is otherwise done in the context of tensor networks by simply growing the bond dimension, although there do exist schemes to improve the accuracy of BP through a preliminary clustering of a network BP15 .
@@ II Belief Propagation
which is also known as the Bethe free energy BP0 .
@@ III Configurations in the BP Basis
with $\mathbb{I}$ as the $d$-dim identity, see also Fig. 1(a). We henceforth refer to $\mathcal{P}_{rs}$ as the projector onto the (BP) ground state of edge $(r,s)$ and $\mathcal{P}_{rs}^{\textrm{C}}$ as the projector onto the excited subspace. It follows that the tensor network $\mathcal{T}$ can be resolved as a sum over $2^{M}$ (BP basis) configurations, formed from all combinations of projecting each edge into either the ground or excited sub-space. We define the degree of a configuration as the number of excited edges that it contains, using $\delta_{x}$ to denote a degree-$x$ configuration and $W(\delta_{x})$ to denote its weight (i.e. the scalar resulting from its contraction). Thus we can recognize $\delta_{0}$ as the BP vacuum; the unique configuration with all edges projected into their BP ground state. It follows that $W(\delta_{0})$ evaluates to the standard BP result of Eq. 4. Our goal is now to characterize the weights of the remaining configurations, which may be thought of as excitations on top of the BP vacuum, such that the high-weight contributions can then be re-introduced into the approximation.
@@
see also the example shown in Fig. 1(c). This result is further simplified by rescaling each tensor $T_{r}$ so that its vacuum state contribution, i.e. $\tilde{T}_{r}$ from Eq. 3, is unity which allows the vacuum component of any configuration to be ignored. This observation implies that one only needs to understand the connected excitations in order to fully characterize a tensor network contraction.
@@ IV.1 Free-energy density
Our first step is to evaluate the weights $W(\delta_{x})$ of the lowest degree excitations $\delta_{x}$; in the calculations performed in this manuscript we evaluate the excitations up to degree $x=14$, see Fig. 2(b-e). Each of these weights is given by contracting a small tensor network formed by capping the external indices (i.e. of tensors within the excitation) with their fixed point messages and projecting the internal indices onto their excited subspace with projectors $\mathcal{P}_{rs}^{\textrm{C}}$ as defined in Eq. 6. However, even once these weights $W(\delta_{x})$ are known, it still remains a non-trivial combinatorial task to compute their impact on the free energy density $f$. Here we shall follow the approximate counting strategy derived in Appendix B.2, which gives the following result for the free energy correction
@@
with $L(\delta_{x})$ as the number of locations per lattice site that excitation can be placed and $S(\delta_{x})$ as the number of hexagonal plaquettes that the excitation occupies. For instance, the $\delta_{6}$ excitation shown in Fig. 1(b) has $L(\delta_{6})=1/2$ (since there is one hexagon per two lattice sites) and $S(\delta_{6})=7$. Given that Eq. 11 contains the unknown $f$ on both sides we cannot solve for $f$ directly. Instead we solve Eq. 11 iteratively by first setting $f=0$ on the r.h.s, computing a new value for $f$, then feeding the new $f$ back into r.h.s and repeating until $f$ is converged.
@@ IV.2 Transfer matrix
We now examine how one can compute the transfer matrix $\textrm{T}_{AB}$ associated to cutting open a single edge from the network $\mathcal{T}=\mbox{$\langle\psi|\psi\rangle$}$ as shown in Fig. 3(a). Such transfer matrices are utilized in determining the optimal truncation of an internal index from the PEPS and thus play an important role in many PEPS optimization algorithms PEPS4 ; PEPS5 ; PEPS6 ; PEPS7 ; PEPS7b ; PEPS7c .
The BP vacuum contribution to $\textrm{T}_{AB}$ is shown in Fig. 3(b) together in Fig. 3(c) with examples of contributions from loop excitations, all of which individually evaluate to a matrix with the same dimensions as $\textrm{T}_{AB}$. The series approximation to $\textrm{T}_{AB}$ is constructed by evaluating each of the loop-excitation networks (up to some desired degree), then adding the contributions together after weighting each by a suppression factor $\exp(S(\delta_{x})f)$ similar to Eq. 11. Note this requires that the free energy density $f$ be evaluated prior to computing $\textrm{T}_{AB}$. Finally, we remark that the excitation terms that do not include either of the tensors adjacent to an open index do not need to be considered explicitly; their contributions, which are proportionate to the vacuum term, are already accounted for through the suppression factors.
@@ V Benchmark results
We now benchmark our approach, beginning with contraction of a PEPS representing the ground state $|\psi\rangle$ of the hexagonal lattice AKLT modelAKLT1 ; AKLT2 in the thermodynamic limit. The AKLT model is particular useful as a testbed, since $|\psi\rangle$ can be represented exactlyPEPS1 with an iPEPS of bond dimension $m=2$. Following the methodology prescribed in Sect. IV we compute the free energy density $f$, transfer matrix $\textrm{T}_{AB}$, and the density matrix $\rho_{AB}$ for loop expansions containing up to $14^{\textrm{th}}$ degree excitations $\delta_{14}$. In all cases we compare against evaluations using a boundary MPS contractionPEPS5 whose dimension $\chi$ has been chosen large enough, $\chi\approx 30$, such that any truncation errors are negligible and the results can be considered as numerically exact. Fig. 4 shows that the addition of the loop corrections improve the accuracy of the BP results dramatically, such that they appear to converge towards the exact results as the degree of the expansion is increased. The expansion to $14^{\textrm{th}}$ degree excitations $\delta_{14}$ gave the free energy density $f$ accurate to $\varepsilon=3\times 10^{-7}$, which represents an improvement of four orders of magnitude over the BP result. That the infinite hexagonal tensor network can be characterised so accurately from only the BP vacuum plus a few of the low-degree loop excitations is remarkable. Note that we also compare against results obtained from the exact contraction of the PEPS on a finite $N\times N$ geometry (with periodic boundaries) for $N=\{6,8,10,12\}$. In all cases we see that an $N^{\textrm{th}}$ degree loop expansion roughly approximates the accuracy of an exact evaluation on a $N\times N$ geometry, providing additional confirmation that the loop expansions are properly accounting for all short-range contributions.
@@
Fig. 5(b-c) presents equivalent results for kagome and square lattices. For iPEPS defined on kagome lattices we perform local manipulations to exactly transform the network into a hexagonal geometry (similar to the form of PEPS described in Ref. PESS1, ) prior to applying the series expansion, see Appendix C for details. The square lattice results were obtained by applying the expansion directly to the square lattice, evaluated up to the $\delta_{8}$ excitations as shown in Appendix C. Similar to the hexagonal lattice, the results from the kagome and square lattices also show robust improvements over the BP approximation. However, we see that the convergence on the square lattice is slower with respect to increasing the degree of expansion; we attribute this to the fact that the quantity of distinct loop configurations at any given degree grows much faster than on the hexagonal lattice.
Figure 4: Results comparing the accuracy of BP and loop expansion up to degree $\delta_{x}$ against exact results for the ground state of the (thermodynamic limit) hexagonal AKLT model. Displayed are errors in (a) the free-energy density $f$, (b) the transfer matrix $\textrm{T}_{AB}$, (c) the two-site density matrix $\rho_{AB}$. Dashed lines show results from an exact contraction of an $N\times N$ system.
@@ VI Discussion
Another direct application of these results could be for PEPS optimization algorithms. Given that a recent workBP13 established the equivalence of BP to the ‘simple update’ strategy for PEPS, the series expansion could significantly enhance the accuracy of the ‘simple update’ strategy while still remaining relatively computationally efficient. For instance, the genus-1 loop excitations on the hexagonal lattice, i.e. the $\delta_{6}$ and $\delta_{10}$ terms in Fig. 2, can be evaluated with $O(m^{6})$ cost for PEPS of bond dimension $m$, and this cost could also potentially be reduced via approximation strategies like sparse diagonalization. Thus the dominant terms in the expansion are still feasible to compute even for PEPS with bond dims in excess of $m=100$, well beyond the computational limits of boundary MPS or corner transfer matrix methods for PEPS contractionsPEPS5 ; PEPS6 ; PEPS7 ; PEPS7b ; PEPS7c ; PEPS8 . A general comparison of the computational efficiency of loop expansions versus MPS-based methods for iPEPS contractions remains beyond the scope of the present manuscript, given that such a comparison would be heavily dependent on the properties of the state under consideration. None-the-less, in Appendix. D we demonstrate that certain (short-range correlated) PEPS states can indeed be contracted more efficiently using the loop expansion than with MPS-based methods.
@@
Most existing tensor network algorithms are built on the twin pillars of tensor contractions and decompositions. The results presented in this manuscript, which prescribe a general procedure for resolving a complicated network into its fundamental components, could represent a third pillar for tensor network algorithms; one that could have interesting synergies with established routines. Exploring the ways that series expansions could compliment or enhance existing tensor network algorithms remains an interesting avenue for future research.
@@ Acknowledgements.
We thank AWS for supporting the quantum computing program. GKC acknowledges support from the US DOE, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (QSA), and a generous gift from AWS. We acknowledge funding provided by the Institute for Quantum Information and Matter, an NSF Physics Frontiers Center (NSF Grant PHY-2317110).
@@ Appendix A Exponential Suppression of High-degree Excitations
The utility of the proposed loop expansion is built on the premise that the weights of excitations (i.e. on top of the BP vacuum) are exponentially diminishing in their degree; this property allows us to accurately approximate a tensor network contraction using only a few low-degree excitations. In this appendix we attempt to justify this premise.
@@
Although our argument assumed, for simplicity, that the transfer matrices with any loop were identical this condition is not strictly necessary; it is only required that the transfer matrices have a gap between their dominant and sub-leading eigenvalues in order to realize (approximate) exponential decay. Finding a general characterisation of the networks that are expected to have a gap in the transfer matrix spectrum associated to any closed loop is likely a difficult task (and certainly beyond the scope of this manuscript). However, given that the loop series expansion is only expected to be viable for networks where the loop corrections are small in comparison to the BP approximation, it follows that the gap condition must hold in this circumstance. Finally we remark that our argument is only valid for excitations that realize genus-1 loops not, for instance, higher genus loops like the $\delta_{11}$ excitations shown in Fig. 2(d). None-the-less, in practice the scaling of Eq. 15 does seem to hold for higher genus loop excitations.